in order to synthesis ferrocene, a student used 6.5 g of FeCl2 x 4H2O and 5.5...

  1. in order to synthesis ferrocene, a student used 6.5 g of FeCl2 x 4H2O and 5.5 mL of cyclopentadiene. the student obtained 2.8 g of ferrocene after sublimation.

Calculate the yield of ferrocene. (Density of cyclopentadiene is 0.786 g/mL)
Answer #1

Balanced equation for the synthesis of ferrocene

FeCl2.4H2O + 2 cyclopentadiene -----> Ferrocene

Number of moles = Mass in gram / Molar mass

Molar mass of FeCl2.4H2O = 198.83 g/mol

Number of moles of FeCl2.4H2O = 6.5 g / 198.83 g/mol = 0.0327 mol

We know mass = Volume * density

Mass of cyclopentadiene = 5.5 ml * 0.786 g/ml = 4.323 g

Molar mass of cyclopentadiene = 66.103 g/mol

Number of moles of ferrocene = 4.323 g / 66.103 g/mol = 0.0654 mol

From reaction, 1.0 mol of FeCl2.4H2O reacts with 2.0 mol of cyclopentadiene produces 1.0 mol of ferrocene then 0.0327 mol of FeCl2.4H2O will react with 2.0 mol * 0.0327 mol / 1.0 mol = 0.0654 mol of cyclopentadiene will produce 0.0327 mol of ferrocene.

Molar mass of ferrocene = 186.04 g/mol

Theoretical yield of ferrocene = 186.04 g/mol * 0.0327 mol = 6.08 g

% yield = Actual yield * 100 / Theoretical yield = 2.8 g * 100 /6.08 g = 46.05 %